OpenStudy (anonymous):

The area of a rectangle is 6 + 3x square feet. Which of the following expressions could represent the lengths of the sides of the rectangle? A. 6 and 3x B. 3 and 2 + x C. 6 and 6 + x D. 3 and 2 + 3x

OpenStudy (anonymous):

Multiply each answer choice together and see which one equals 6 + 3x 6 times 3x = 3 times (2 + x) = 6 times (6 + x) = 3 times (2 + 3x) =